A farmer has 1,080 acres of land on which he grows corn, wheat, and soybeans. It costs $45 per acre to grow corn, $60 to grow wheat, and $50 to grow soybeans. Because of market demand the farmer will grow twice as many acres of wheat as of corn. He has allocated $57,300 for the cost of growing his crops. How many acres of each crop should he plant?
Found 2 solutions by ankor@dixie-net.com, ikleyn:
Answer by ankor@dixie-net.com(22740)

A farmer has 1,080 acres of land on which he grows corn, wheat, and soybeans.
It costs $45 per acre to grow corn, $60 to grow wheat, and $50 to grow soybeans.
Because of market demand the farmer will grow twice as many acres of wheat as of corn.
He has allocated $57,300 for the cost of growing his crops.
How many acres of each crop should he plant?
:
let c = acres in corn
let w = acres in wheat
let s = acres in soy bean
:
write an equation for each statement
:
"A farmer has 1,080 acres of land on which he grows corn, wheat, and soybeans."
c + w + s = 1080
" It costs $45 per acre to grow corn, $60 to grow wheat, and $50 to grow soybeans.".."He has allocated $57,300 for the cost"
45c + 60w + 50s = 57300
"Because of market demand the farmer will grow twice as many acres of wheat as of corn."
w = 2c
:
replace w with 2c in the first two equations
c + 2c + s = 1080
3c + s = 1080
and
45c + 60(2c) + 50s = 57300
45c + 120c + 50s = 57300
165c + 50s = 57300
simplify, divide equation by 50
3.3c + s = 1146
use elimination on these two equation
3.3c + s = 1146
3c + s = 1080
--------------------subtraction eliminates s, find c
.3c + 0 = 66
c = 66/.3
c = 220 acres in corn
then
w = 2(220)
w = 440 acres in wheat
then
1080 - 220 - 440 = 420 acres in soybeans
:
:
Check this in the statement:
" It costs $45 per acre to grow corn, $60 to grow wheat, and $50 to grow soybeans.".."He has allocated $57,300 for the cost"
45(220) + 60(440) + 50(420) =
9900 + 26400 + 21000 = 57300

Answer by ikleyn(52788)

.

            There is much simpler way to solve the problem,  using  one  SINGLE  equation in one unknown
            INSTEAD  of system of  3  equations in  3  unknowns.


Let x be the area for corn, in acres.

Then the area for wheat is 2x acres.

Hence, the area for soybeans is  (1080 - x - 2x) = 1080-3x acres.


Now it is easy  to write the "money" equation


    45*x + 60*(2x) + 50*(1080-3x) = 57300.


Simplify and solve for x


    45x + 120x + 50*1080 - 150x = 57300

    15x = 57300 - 50*1080  = 3300

       x                   = 3300/15 = 220.


ANSWER.  220 acres for corn;  2*220 = 440 acres for wheat and the rest,  1080-220-440 = 420 acres for soybeans.

Solved and completed.

------------

The whole problem is designed, intended and expected to be solved exactly  IN  THIS  WAY.

It was given to you  WITH  THE  ONLY  GOAL :   in order for you learned from my post how to solve it in the  RIGHT  WAY.